OpenStudy (christos):

Can someone help me find the exact value of - sec10* /csc80*

OpenStudy (christos):

please help me please

OpenStudy (anonymous):

do you understand the concept of secant and cosecant first of all?

OpenStudy (christos):

Yes, there might be some gaps in between but I can get the bigger picture :D

OpenStudy (christos):

@all

OpenStudy (christos):

@everyone

OpenStudy (anonymous):

hmm how about the unit circle?

OpenStudy (christos):

No my friend, I donnto do graph just equations and stuff like that :)

OpenStudy (christos):

I just solve broblems without graph

OpenStudy (anonymous):

okay so what is the sine 10* and cosine of 80*

hartnn (hartnn):

sec (90-y) = csc (y) put y=80 here, what u get ?

OpenStudy (christos):

I stil dont understand how To solve this problem!!!

hartnn (hartnn):

:O \(\sin (90-x)=\cos x \\ \cos(90-x)=\sin x\\ \tan (90-x)=\cot x \\ \cot(90-x)=\tan x\\ \sec(90-x)= \csc x \\ \csc(90-x)=\sec x \) so, using \(\sec(90-x)= \csc x\) put x= 80 , what u get ??

hartnn (hartnn):

why is it so difficult to put x=80 in that ?